# A tuneable frequency comb via dual-beam laser-solid harmonic generation

### v2

@@ Abstract
A high-power laser pulse at normal incidence onto a plane solid target will generate odd harmonics of its frequency. However, the spacing of the harmonic lines in this configuration is fixed. Here, we study harmonic generation using two laser beams incident on a plane target at small, opposite angles to the target normal, via particle-in-cell simulations. When looking at the harmonic radiation in a specific direction via a narrow slit or pinhole, we select an angle-dependent subset of the harmonic spectrum. This way, we obtain a harmonic frequency comb that we control via the observation angle and the input laser frequency. The divergence of the harmonic radiation will be reduced by using wider laser spots, thus increasing the efficacy of the scheme. We will discuss extensions to this scheme, such as using beams with unequal frequencies, a slight tilt of the target, or employing more than two beams.
@@ I Introduction
HHG with non-collinear laser pulses has been studied in the past hickstein ; fomichev ; bertrand ; heyl14 ; gariepy ; zli ; negro ; ellis . In most cases, the laser beams had different wave lengths, e.g. a fundamental at 800 nm and its second harmonic at 400 nm. While a 2-D $(\omega,k_{\perp})$ spectrum is usually shown, in most cases the angular diffraction pattern is studied for a given harmonic order. In our work, we aim to start from a given diffraction angle, and study the subset of harmonic orders that is diffracted in that specific angle.
Divergence of a “necklace beam”: the region where the harmonics are generated has a ring shape, so its radius $R_{0}$ will be roughly the same regardless of the harmonic order. The Rayleigh length of harmonic $q\omega_{0}$ is then given by $z_{q}=k_{q}R_{0}^{2}/2=qk_{0}R_{0}^{2}/2=qz_{0}$, explaining why Rego et al. rego22 find a decreasing divergence for. increasing harmonic order. In our work, we use driving pulses with super-Gaussian envelopes to ensure that the effective radius of the harmonic-generating region does not decrease too much for increasing harmonic order.
@@ II Theory
For the original necklace beam, the laser configuration is given by the four points ±$\pm(1,\ell_{A,B})$ in $(\omega/\sigma,\ell/\sigma)$ space. The harmonic progression then has two degrees of freedom and is given by:
@@
For a scenario with two identical s-polarised LP beams crossing at a shallow angle, as used by e.g. Hickstein et al. hickstein , the harmonic progression reads:
@@
One can consider a generalisation to $ck_{\perp}/\omega=\tan(15^{\circ})N_{\perp}/N_{\parallel}$ with $\gcd(N_{\perp},N_{\parallel})=1$. However, this does not yield different spectra. Since $q=2m+1$ must be an odd integer, one does not get a different frequency comb for $N_{\perp}\not=1$, e.g. $ck_{\perp}/(\omega/\tan(15^{\circ}))=3/7$ instead of $1/7$. But in experiments, it may help to have multiple available angles to obtain a specific comb.
@@ III Simulation setup
We carried out 2-D particle-in-cell simulations using the Epoch code epoch1 ; epoch2 to model the interaction of two laser beams hitting a flat target at near-normal incidence. Numerical parameters were: a box size of $16\times 64\ \mu\mathrm{m}^{2}$, a grid size of $4096\times 8192$ cells, 10 particles per cell, and a total simulation time of 250 fs. Physical parameters were: a laser wave length of 1 $\mu$m, a laser peak intensity of $3\cdot 10^{19}$ W/cm2, a beam waist of 16 $\mu$m, a transverse pulse envelope that is either Gaussian ($\exp(-r^{2})$) or super-Gaussian ($\exp(-r^{16})$), a Gaussian longitudinal envelope with a duration of 50 fs, a plasma slab with a maximum density of $8n_{\mathrm{cr}}$ and a pre-plasma layer extending 1 $\mu$m in front of the slab. The laser beams are identical with linear s-polarisation and impact the same spot at angles of ±$\pm 15$ degrees with respect to the target normal.
@@
We need a fairly wide spot diameter: $k_{0}R\tan(15^{\circ})>1$. Otherwise: (i) not enough “grating” periods within the spot, so harmonic amplitude becomes poor, (ii) large divergence angle in the harmonics, so harmonics from different emission angles overlap and selecting a specific viewing angle becomes difficult or impossible, (iii) the divergence angle for higher harmonics will be compromised rego22 : a given harmonic needs a minimum intensity to be generated “non-perturbatively”, so the spot area where this intensity is reached should be large enough to ensure a narrow divergence angle in the far field. In Ref. rego22 the pump laser spot is always a ring, so does not really get smaller for a higher intensity threshold, but in our case the spot diameter will get smaller for a higher intensity threshold, so one start with a decent spot diameter and laser power in the first place. The effects of this can also be seen in the poor harmonic divergence obtained for Gaussian driving beams, and the much improved divergence obtained for super-Gaussian beams (see below).
IV Simulation results
@@
Figure 1: Comparison of the quality of the spectral peaks generated by either (a,b) a Gaussian beam or (c,d) a super-Gaussian beam. For the Gaussian beam, the divergence of the harmonic peaks is too large to be practical, see frame (a), while for the super-Gaussian beam, the divergence has much improved, see (c). This can be explained from the contours of the beam envelope: for a Gaussian beam, the radius of the contours decreases quickly for increasing energy, see (b), while for the super-Gaussian, the contour radius remains much larger for increasing energy. Thus, in the case of a super-Gaussian beam, the harmonics are generated from a much larger “effective spot diameter” than for a Gaussian beam, leading to improved harmonic divergence.
@@
Figure 2: The $(\omega/\sigma,k_{\perp}/\sigma)$ spectrum of the reflected harmonic radiation. The positions of the peaks are as predicted by Eq. (1). Line-outs are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1,3,5$, as indicated in the plot.
(a) Refer to caption (b) Refer to caption
(c) Refer to caption (d) Refer to caption
(e) Refer to caption (f) Refer to caption
Figure 3: The “signed frequency” $\omega/\sigma$ spectrum corresponding to the line-outs from Figure 2. (a,c,e): results for Gaussian driving beams $\exp(-r^{2})$; (b,d,f): results for super-Gaussian beams $\exp(-r^{16})$. Cuts are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1$ (a,b), $N=3$ (c,d) and $N=5$ (e,f). In each case, spectral peaks are found at $\omega/\sigma=(2n+1)N$, $n\in\mathbb{Z}$. The spectra for siuper-Gaussian driving beams are much clearer than those for Gaussian beams, as discussed in the text. Each image displays a frequency comb with a different line separation, while all images are taken from the results of the same simulation. This proves that a multitude of frequency combs can be obtained from the same laser-target configuration.
@@ V Extensions
    When the harmonic progression has more than 2 dimensions, e.g. 3-D, then one can either choose a 2-D subset on a plane or a 1-D subset on a line. This offers an extra degree of freedom in selecting frequency combs and thus deserves further exploration.
@@ VI Conclusions
We have studied HHG in the reflection of two identical high-power laser beams off a plane solid target, where the laser beams have small, opposite angles of incidence with respect to the target normal. The reflected harmonic light does not just contain many frequencies, but also many values for the perpendicular wave number $k_{\perp}$. For a given oblique observation angle $\alpha$, only those harmonic frequencies that have a component for which $\tan\alpha=k_{\perp}/k_{\parallel}$ will be seen. By carefully selecting harmonic light in only a specific direction, a harmonic frequency comb is found where the spacing of the comb depends on the chosen direction. Multiple frequency combs can be obtained from the same simulation, and may thus also be obtained from the same experiment. This versatile approach admits multiple extensions, several of which are listed in this paper.

### v3

@@ Abstract
A high-power laser pulse at normal incidence onto a plane solid target will generate odd harmonics of its frequency. However, the spacing of the harmonic lines in this configuration is fixed. Here, we study harmonic generation using two laser beams incident on a plane target at small, opposite angles to the target normal, via particle-in-cell simulations. When looking at the harmonic radiation in a specific direction via a narrow slit or pinhole, we select an angle-dependent subset of the harmonic spectrum. This way, we obtain a harmonic frequency comb that we control via the observation angle and the input laser frequency. The spectra for s- and p-polarised harmonics are studied separately, as they offer different frequency combs. The divergence of the harmonic radiation will be reduced by using wider laser spots, thus increasing the efficacy of the scheme. We will discuss extensions to this scheme, such as using beams with unequal frequencies, a slight tilt of the target, or employing more than two beams.
@@ I Introduction
HHG with non-collinear laser pulses has been studied in the past hickstein ; fomichev ; bertrand ; heyl14 ; gariepy ; zli ; negro ; ellis . In most cases, the laser beams had different wave lengths, e.g. a fundamental at 800 nm and its second harmonic at 400 nm. While a 2-D $(\omega,k_{\perp})$ spectrum is usually shown, in most cases the angular diffraction pattern is studied for a given harmonic order. In our work, we aim to start from a given diffraction angle, and study the subset of harmonic orders that is diffracted in that specific angle. We also study the harmonic spectra for s- and p-polarised harmonics separately, as these offer different harmonic progressions and thus different options for frequency combs.
The divergence of the harmonic modes deserves special attention. For a Gaussian driving beam, the effective radius of the part of the envelope that generates a given harmonic, i.e. where the pulse intensity is above a given threshold, decreases rapidly with increasing harmonic order. As a result, higher order harmonics are generated by an ever smaller part of the pump beam(s), so their divergence increases with harmonic order. For a “necklace beam”: the region where the harmonics are generated has a ring shape, so its radius $R_{0}$ will be roughly the same regardless of the harmonic order. The Rayleigh length of harmonic $q\omega_{0}$ is then given by $z_{q}=k_{q}R_{0}^{2}/2=qk_{0}R_{0}^{2}/2=qz_{0}$, explaining why Rego et al. rego22 find a decreasing divergence for. increasing harmonic order. In our work, we use driving pulses with super-Gaussian envelopes to ensure that the effective radius of the harmonic-generating region does not decrease too much for increasing harmonic order.
@@ II Theory
For the original necklace beam rego22 , the laser configuration is given by the four points ±$\pm(1,\ell_{A,B})$ in $(\omega/\sigma,\ell/\sigma)$ space. The harmonic progression then has two degrees of freedom and is given by:
@@
For a scenario with two identical s-polarised LP beams crossing at a shallow angle $2\alpha_{0}$ (we will use $\alpha_{0}=15^{\circ}$ in our simulations below), as used by e.g. Hickstein et al. hickstein , the vector potential $\mathbf{A}_{0}\propto\mathbf{e}_{s}$ of the driving laser beams is given by the four points (±$(\pm 1,\pm k_{0\perp})$ in $(\omega/\sigma,k_{\perp}/\sigma)$ space, with $k_{0\perp}=k_{0}\sin\alpha_{0}$. The lowest order term driving the s-polarised harmonics is then $A_{0}^{2}\mathbf{A}_{0}$. The harmonic progression for the s-polarised light has two degrees of freedom and is given by:
@@
One can consider a generalisation to $ck_{\perp}/\omega=\tan(\alpha_{0})N_{\perp}/N_{\parallel}$ with $\gcd(N_{\perp},N_{\parallel})=1$. However, this does not yield different spectra. Since $q=2m+1$ must be an odd integer, one does not get a different frequency comb for $N_{\perp}\not=1$, e.g. $ck_{\perp}/(\omega/\tan(\alpha_{0}))=3/7$ instead of $1/7$. But in experiments, it may help to have multiple available angles to obtain a specific comb.
In addition to the s-polarised harmonics, we note that the work by Lichters et al. lichters also predicts p-polarised harmonics at even multiples of the base frequency and wave vector, even for purely s-polarised pump beams. These harmonics are emitted by the “DC mode” $\mathbf{A}_{\mathrm{DC}}=\tan(\alpha_{0})\mathbf{e}_{p}$ via the nonlinear term $A_{0}^{2}\mathbf{A}_{\mathrm{DC}}$. The harmonic progression for the p-polarised light also has two degrees of freedom and is given by
$(\omega/\sigma,k_{\perp}/\sigma)_{mn}=m(2,0)+n(0,2).$ (2)
For these harmonics, the emission half angle of any harmonic mode is given by $\tan\alpha_{mn}=(2n)k_{\perp 0}/[(2m)k_{\parallel 0}]=(n/m)\tan\alpha_{0}$. For e.g. a viewing angle $\alpha_{v}$ given by $\tan\alpha_{v}=k_{\perp 0}/(Nk_{\parallel 0})$, with $N$ any integer, we preferentially select modes $q\omega_{0}$ with $q=2m=(2n)N$, $n\in\mathbb{N}$. We note two differences with the s-polarised frequency combs give above: (i) any value for $N$ is now possible, not just odd ones, and (ii) the frequency of the lowest harmonic is now the same as the frequency step size, rather than half the step size.
@@ III Simulation setup
We carried out 2-D particle-in-cell simulations using the Epoch code epoch1 ; epoch2 to model the interaction of two laser beams hitting a flat target at near-normal incidence. Numerical parameters were: a box size of $16\times 64\ \mu\mathrm{m}^{2}$, a grid size of $4096\times 8192$ cells, 10 particles per cell, and a total simulation time of 250 fs. Physical parameters were: a laser wave length of 1 $\mu$m, a laser peak intensity of $3\cdot 10^{19}$ W/cm2, a beam waist of 16 $\mu$m, a transverse pulse envelope that is either Gaussian ($\exp(-r^{2})$) or super-Gaussian ($\exp(-r^{16})$), a Gaussian longitudinal envelope with a duration of 50 fs, a plasma slab with a maximum density of $8n_{\mathrm{cr}}$ and a pre-plasma layer extending 1 $\mu$m in front of the slab. The laser beams are identical with linear s-polarisation and impact the same spot at angles of ±$\pm\alpha_{0}=\pm 15^{\circ}$ with respect to the target normal.
@@
We need a fairly wide spot diameter: $k_{0}R\tan(15^{\circ})>1$. Otherwise: (i) there will not be enough “grating” periods within the spot, so the harmonic amplitude becomes poor, (ii) the harmonics will have a large divergence angle, so harmonics from different emission angles overlap and selecting a specific viewing angle becomes difficult or impossible, (iii) the divergence angle for higher harmonics will be compromised rego22 : a given harmonic needs a minimum intensity to be generated “non-perturbatively”, so the spot area where this intensity is reached should be large enough to ensure a narrow divergence angle in the far field. In Ref. rego22 the pump laser spot is always a ring, so does not really get smaller for a higher intensity threshold, but in our case the spot diameter will get smaller for a higher intensity threshold, so one start with a decent spot diameter and laser power in the first place. The effects of this can also be seen in the poor harmonic divergence obtained for Gaussian driving beams, and the much improved divergence obtained for super-Gaussian beams (see below).
@@
Figure 1: Comparison of the quality of the s-polarised spectral peaks generated by either (a,b) two Gaussian beams or (c,d) two super-Gaussian beams. For the Gaussian beams, the divergence of the harmonic peaks is too large to be practical, see frame (a), while for the super-Gaussian beams, the divergence has much improved, see (c). This can be explained from the contours of the beam envelope: for a Gaussian intensity envelope, the radius of the contours decreases quickly for increasing energy, see (b), while for the super-Gaussian envelope, the contour radius remains much larger for increasing energy. Thus, in the case of a super-Gaussian beam, the harmonics are generated from a much larger “effective spot diameter” than for a Gaussian beam, leading to improved harmonic divergence.
IV Simulation results
IV.1 S-polarised harmonics
The s-polarised harmonics, as given by Eq. (1), all have frequencies and transverse wave numbers that are odd multiples of those of the pump beams. These harmonics do not require the presence of any DC mode to be generated lichters ; trines24 , only the presence of a cubic nonlinear term in the wave equation. This sequence of harmonics can also be found in experiments on HHG in noble gases hickstein ; rego22 , unlike the p-polarised harmonics below, which are more particular to our configuration of two laser pulses hitting a solid surface. The results for this case are shown in Figures 1-3 and summarised below.
Figure 2: The $(\omega/\sigma,k_{\perp}/\sigma)$ spectrum of the reflected s-polarised harmonic radiation for the super-Gaussian beams. The positions of the peaks are as predicted by Eq. (1). Line-outs are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1,3,5$, as indicated in the plot (smaller slope means higher $N$).
(a) Refer to caption (b) Refer to caption
(c) Refer to caption (d) Refer to caption
(e) Refer to caption (f) Refer to caption
Figure 3: The “signed frequency” $\omega/\sigma$ spectrum of the s-polarised radiation corresponding to the line-outs from Figure 2. (a,c,e): results for Gaussian driving beams $\exp(-r^{2})$; (b,d,f): results for super-Gaussian beams $\exp(-r^{16})$. Cuts are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1$ (a,b), $N=3$ (c,d) and $N=5$ (e,f). In each case, spectral peaks are found at $\omega/\sigma=(2n+1)N$, $n\in\mathbb{Z}$. The spectra for siuper-Gaussian driving beams are much clearer than those for Gaussian beams, as discussed in the text. Each image displays a frequency comb with a different line separation, while all images are taken from the results of the same simulation. This proves that a multitude of frequency combs can be obtained from the same laser-target configuration.
@@ IV.1 S-polarised harmonics
IV.2 P-polarised harmonics
From the work by Lichters et al. lichters , it follows that an s-polarised pump beam, obliquely incident onto a flat surface, will generate even harmonics with p-polarisation. This can be explained as follows. Let $\mathbf{e}_{s}$ and $\mathbf{e}_{p}$ denote the unit vectors for s- and p-polarisation, with $\mathbf{e}_{s}\cdot\mathbf{e}_{p}=0$. Then the pump vector potential is $\mathbf{A}_{0}\propto\mathbf{e}_{s}$, while our specific laser-target configuration results in a DC mode $\mathbf{A}_{\mathrm{DC}}\propto(m_{e}c/e)\tan(15^{\circ})\mathbf{e}_{p}$, and a nonlinear term $\propto A_{0}^{2}\mathbf{A}_{\mathrm{DC}}$ in the wave equation, emitting even harmonics with p-polarisation. As discussed above, the harmonic spectrum is given by $(\omega/\sigma,k_{\perp}/\sigma)_{mn}=(2m,2n)$ for $m,n\in\mathbb{Z}$, and for a viewing angle $\alpha_{v}$ given by $\tan\alpha_{v}=k_{\perp 0}/(Nk_{\parallel 0})=tan(15^{\circ})/N$, with $N$ any integer, we preferentially select modes $q\omega_{0}$ with $q=2m=(2n)N$, $n\in\mathbb{N}$. Simulation results for the p-polarised light are displayed in Figures 4 (full $(\omega/\sigma,k_{\perp}/\sigma)$ spectrum) and 5 (line-outs at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1,2,3$). We note that these results were obtained from the simulation with the super-Gaussian driving beams discussed above (no separate simulation needed).
Figure 4: The $(\omega/\sigma,k_{\perp}/\sigma)$ spectrum of the reflected p-polarised harmonic radiation from the super-Gaussian simulation. The positions of the peaks are as predicted by Eq. (2). Line-outs are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1,2,3$, as depicted in the plot (smaller slope means higher $N$). The p-polarised spectrum in this figure is taken from the same simulation as the s-polarised spectrum of Figure 3, once again emphasising that a multitude of frequency combs can be obtained from the same laser-target configuration.
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 5: Results for the p-polarised light from the super-Gaussian simulation. Line-outs are taken at $ck_{\perp}/\omega=\tan(15^{\circ})/N$ for $N=1,2,3$. The “signed frequency” $\omega/\sigma$ spectrum of the p-polarised radiation corresponding to the line-outs taken from Figure 4 for (a) $N=1$, (b) $N=2$ and (c) $N=3$. In each case, we see that the lowest spectral peaks are at $\omega/(\sigma\omega_{0})=\pm N$. While the resolution of our simulation is not sufficient to see a multitude of spectral peaks in the line-outs, the spectrum in Figure 4 proves that the underlying principle is sound.
@@ V Extensions
    When the harmonic progression has more than 2 dimensions, e.g. 3-D, then one can either choose a 2-D subset on a plane or a 1-D subset on a line. This offers an extra degree of freedom in selecting frequency combs and thus deserves further exploration in a separate publication.
@@ VI Conclusions
We have studied HHG in the reflection of two identical high-power laser beams off a plane solid target, where the laser beams have small, opposite angles of incidence with respect to the target normal. The reflected harmonic light does not just contain many frequencies, but also many values for the perpendicular wave number $k_{\perp}$. For a given oblique observation angle $\alpha$, only those harmonic frequencies that have a component for which $\tan\alpha=k_{\perp}/k_{\parallel}$ will be seen. By carefully selecting harmonic light in only a specific direction, a harmonic frequency comb is found where the spacing of the comb depends on the chosen direction. Multiple frequency combs can be obtained from the same simulation, and may thus also be obtained from the same experiment. Also, the s- and p-polarised components of the harmonic light provide different harmonic progressions and thus separate sets of frequency combs, with either odd or even multiples of the pump frequency $\omega_{0}$. This versatile approach admits multiple extensions, several of which are listed in this paper.
@@
Acknowledgements
One of the authors (RT) would like to thank Eva Los and Chris Murphy for useful discussions. Simulations were performed on the Scarf computer cluster at the STFC Rutherford Appleton Laboratory.